OpenStudy (anonymous):

The function f(x) is shown below: f(x) = (3)x - 2 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? f(x) = (3)x + 11 f(x) = (3)x + 13 f(x) = (3)x -15 f(x) = (3)x - 13

OpenStudy (solomonzelman):

\(\large\color{ blue }{\large {\bbox[5pt, lightgreen ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline~~~~~~~~~~~~~~~~~~~~~~~~~~~\textbf{Shifts}~~~~~~~~~~~~~~~~~~~~~~~~~~~&~\bf{c~~~units~~~~} \\ \hline \\f(x)= ∛x ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{red }{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= \sqrt[3]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= ∛x \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= ∛x ~~~ ⇒ ~~~ f(x)= ∛x \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

in your case, what do you think you should do to shift it 13 units to the left ?